# Structure growth in $f(Q)$ cosmology

### v1

@@ 3 Cosmological perturbations in $f(Q)$ gravity
where the frictional force, restoring force, and source force are expressed by $A$, $B$, and $C$, respectively, and the separation of variables takes the form $X=\sum_{k}X^{k}(t)Q^{k}(x),\hskip 2.84544pt{\rm and}\hskip 2.84544ptY=\sum_{k}Y^{k}(t)Q^{k}(x)\;,$ where $k$ is the wave number and $Q^{k}(x)$ is the eigenfunction of the covariantly defined Laplace-Beltrami operator in (almost) FLRW space-times, $\nabla^{2}Q^{k}(x)=-\frac{k^{2}}{a^{2}}Q^{k}(x).$ Then, the first-order evolution equations (3.13) - (3.16) become:
$\dot{\mathcal{D}}^{k}_{m}=-(1+w)Z^{k}+w\theta\mathcal{D}^{k}_{m}\;,$ (3.18)
$\dot{Z}^{k}=-\Big{[}\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\Big{]}Z^{k}+\Bigg{[}\frac{w\theta^{2}}{3(1+w)}-\frac{(1+3w)}{2(1+w)}\rho_{m}$
$-\frac{wF}{2(1+w)}-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}+\frac{k^{2}}{a^{2}}\frac{w}{1+w}\Bigg{]}\mathcal{D}^{k}_{m}$
$+\Bigg{[}\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{]}\mathcal{W}^{k}-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}^{k}$ (3.19)
$\dot{\mathcal{W}}^{k}=\mathcal{L}^{k}-\frac{w\dot{Q}}{1+w}\mathcal{D}^{k}_{m}\;,$ (3.20)
˙$\dot{\mathcal{L}}^{k}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}^{k}-\frac{w\ddot{Q}}{1+w}\mathcal{D}^{k}_{m}\;.$ (3.21)
@@
Consequently, the second-order evolution equations (3.18) - (3.21) with respect to time derivative yield as
$\ddot{\mathcal{D}}_{m}^{k}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}+\dot{Q}F^{\prime\prime}-w\theta\right]\dot{\mathcal{D}}_{m}^{k}-\left[-wF-\theta\dot{Q}F^{\prime\prime}w+\frac{\theta\dot{Q}F^{\prime}w}{2Q}\right.$
$\left.-\frac{(1+3w)\rho_{m}}{2}(1-w)+\frac{k^{2}}{a^{2}}w\right]\mathcal{D}_{m}^{k}-\left[\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}\right.$
$\left.-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right](1+w)\mathcal{W}^{k}+\left[\frac{\theta F^{\prime}}{2Q}-\theta F^{\prime\prime}\right](1+w)\dot{\mathcal{W}}^{k}\,,$ (3.22)
$\ddot{\mathcal{W}}^{k}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}^{k}-\frac{2w\ddot{Q}}{1+w}\mathcal{D}_{m}-\frac{w\dot{Q}}{1+w}\dot{\mathcal{D}}_{m}^{k}\,.$ (3.23)
3.1 Density contrast
@@
$\dot{Y}=-(1+z)HY^{\prime}\;,\qquad\ddot{Y}=(1+z)H^{2}Y^{\prime}+(1+z)^{2}H^{2}Y^{\prime\prime}+(1+z)^{2}H^{\prime}HY^{\prime}\;,$ (3.24)
@@
to present the growth density fluctuations through cosmological redshift. Thus, by resorting to expressions Eqs. (3.24) and assuming the power law model $F(Q)=\alpha Q^{n}$, the second-order evolution equations Eqs. (3.22) - (3.23) read as follows:
$(1+z)^{2}{\mathcal{D}}^{\prime\prime}_{m}=(1+z)\bigg{\{}1-(1+z)\frac{H^{\prime}}{H}+\frac{H^{\prime}}{H}n\left({1-\Omega_{m}}\right)E^{2n-2}\bigg{\}}{\mathcal{D}}^{\prime}_{m}$
$+\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\mathcal{D}_{m}-\frac{1}{H^{2}}\Bigg{\{}\frac{(1-\Omega_{m})}{(1-2n)}E^{2n-2}\bigg{\{}\frac{n^{2}}{2}-n$ (3.25)
$-(1+z)\frac{{H^{\prime}}}{H}n\bigg{(}n^{2}-4n+\frac{7}{2}\bigg{)}\bigg{\}}\Bigg{\}}\mathcal{W}-(1+z)\left[\frac{n}{2H^{2}}\left({(1-\Omega_{m})}E^{2n-2}\right)\right]{\mathcal{W}}^{\prime}\;,$
$\left(1+z\right)^{2}\mathcal{W}^{\prime\prime}=16\frac{5n(\Omega_{m}+3)-\Omega_{m}-9}{6(2n-1)(n^{2}\Omega_{m}-n^{2}+2n-1)}\mathcal{W}\;.$ (3.26)
@@
As we can see, Eqs. (3.22) - (3.23) are non-linear evolution equations in which no approximation has been made. To convert such a set of equations into a closed second-order equation, many works [20, 19, 32, 33] have considered the quasi-static approximation. For simplicity, we applied the quasi-static approximation, where the first and second order of the time derivative of nonmetric density fluctuations is assumed to be approximately zero ($\dot{\mathcal{W}}=\ddot{\mathcal{W}}\approx 0$). Once the quasi-static approximation is made Eqs. (3.22) - (3.23) reduced to the closed system equation and it becomes
$\ddot{\mathcal{D}}_{m}^{k}=-\Bigg{\{}\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}+\dot{Q}F^{\prime\prime}-w\theta+\Bigg{(}\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}$
@@
$+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{)}\left(\frac{w\dot{Q}^{2}}{\dddot{Q}}\right)\Bigg{\}}\dot{\mathcal{D}}_{m}^{k}-\Bigg{\{}\frac{\theta\dot{Q}F^{\prime}w}{2Q}-wF-\theta\dot{Q}F^{\prime\prime}w-\frac{(1+3w)\rho_{m}}{2}(1-w)$
@@
$+\frac{k^{2}}{a^{2}}w+\Bigg{(}\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{)}\left(\frac{2w\dot{Q}\ddot{Q}}{\dddot{Q}}\right)\Bigg{\}}\mathcal{D}^{k}_{m}\,.$ (3.27)
Once redshift-space transformation is implemented  [^8] For the case of $n=0$, Eqs.(3.27) is reduced to the usual $\Lambda$CDM equation for the density contrast as $\ddot{\mathcal{D}}_{m}^{k}=-\left(\frac{2}{3}-w\right)\theta\dot{\mathcal{D}}_{m}^{k}+\frac{1}{3}\theta^{2}\left[2w(1-\Omega_{m})+\frac{\Omega_{m}(1+3w)}{2}(1-w)-\frac{k^{2}}{a^{2}}w\right]\mathcal{D}_{m}^{k}\;.$ (3.28) , we find that Eq. (3.27) for dust ($\omega=0$) yields
$(1+z)^{2}{\mathcal{D}}^{\prime\prime}_{m}$ $=$ $(1+z)\left[1-(1+z)\frac{H^{\prime}}{H}+\,\frac{H^{\prime}}{H}\left(1-\Omega_{m}\right)nE^{2n-2}\right]{\mathcal{D}}^{\prime}_{m}$ (3.29)
$+$ $\,\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\,\mathcal{D}_{m}\;.$
@@
To solve the density contrast numerically resorting the evolution equations (3.25)-(3.26), we use the initial values of the system. Due to gravitational instability, the variation in the temperature of the CMB is detected to be observational on the order of $10^{-5}$ [34] at $z_{in}\approx 1089$, and we consider it as an initial condition. This small gravitational instability is the seed for the formation of large-scale structures. Hence, we define the normalized energy density contrast for the matter fluid as
${\mathcal{\delta}}(z)=\frac{\mathcal{D}_{m}(z)}{\mathcal{D}_{m}{(z_{in})}}\;,$ (3.30)
@@
where the subscript $in$ refers to the initial value of $\mathcal{D}_{m}(z)$ at the given initial redshift $z_{in}$ [34]. As stated previously, we consider two different situations in this section: the first illustration does not assume any (quasi-static) approximation (dubbed the full system in the following), whereas the second situation relies on the quasi-static approximation to determine the evolution of the density contrast $\delta_{m}(z)$ in the matter-dominated universe.
3.2 The numerical results of $\delta(z)$
We first consider the evolution equation without any approximation as presented in Eqs. (3.25) - (3.26), and show the results for the density contrast in the left outer panel of Fig. 2. using the best-fit values of $n$. Second, we have studied the growth of density fluctuations resorting to the quasi-static approximation as per Eq. (3.29) whose results are presented in the right outer panel of Fig. 2 for the matter-dominated universe. Meanwhile, we have chosen the best fit values of $\Omega_{m}=0.259$ and $n=0.09^{+0.064}_{-0.064}$ for illustrative purposes from OHD + SNIa as presented in Table 1 to represent the density contrast $\delta(z)$ for the model $f(Q)$ and $\Omega_{m}=0.263$ for $\Lambda$CDM as well. The numerical values of the matter density fluctuations $\delta(z)$ turned out to be very sensitive to the constrained values of $\Omega_{m}$ and the exponent $n$ in both scenarios (full and quasi-static).
@@ 3.2 The numerical results of $\delta(z)$
Figure 2: The outer left panel displays growth density fluctuations, $\delta(z)$ when resolving the whole system (3.25)-(3.26), while the inner left panel illustrates the relative difference, $\zeta(z)$, between the $\Lambda$CDM and $f(Q)$-power-law models across the cosmological redshift. The outer right panel depicts growth density fluctuations $\delta(z)$ under the quasi-static approximation given by (3.29), and the inner right panel showcases the relative difference $\zeta(z)$, between the $\Lambda$CDM and $f(Q)$ power-law models. These plots correspond to values of $n=0.09\pm 0.64={0.026,0.09,0.134}$ and $\Omega_{m}=0.259$, which are obtained from the best-fit parameters derived using OHD + SNIa data (see Table 1).
@@
We have also presented the relative difference of $\delta(z)$ between $\Lambda$CDM and $f(Q)$ model by defining the dimensionless parameter $\zeta(z)$ as
$\zeta(z)=\bigg{|}\frac{\delta^{\rm\Lambda CDM}(z)-\delta^{f(Q)}_{\rm{full,QSA}}(z)}{\delta^{\rm\Lambda CDM}(z)}\bigg{|}\;\;,$ (3.31)
@@
Figure 3: The relative difference between the full system and the quasi-static resolutions for $f(Q)$-gravity model for the values of $n=0.09\pm 0.64={0.026,0.09,0.134}$ and $\Omega_{m}=0.263$ which have been taken from the best-fit values obtained using OHD+SNIa data, see Table 1.
@@
$\eta(z)=\bigg{|}\frac{\delta^{f(Q)}_{\rm{QSA}}(z)-\delta^{f(Q)}_{\rm{full}}(z)}{\delta^{f(Q)}_{\rm{full}}(z)}\bigg{|}\;,$ (3.32)
and presented its numerical finding in Fig. 3. This definition serves as an indicator of the competitiveness of quasi-static with respect to the full system resolution. From Fig. 3, we conclude that the quasi-static approximation represents better the evolution of the density contrast for values of $n$ closer to zero ($n=0$ corresponding to $\Lambda$CDM). Based on this result, we highlight that the quasi-static approximation, which has been widely taken into consideration in other modified theories of gravity [17, 19, 35, 36, 37] may be used to represent with a competitive - 84.89% at $n=0.026$, 82.1% at $n=0.09$, and 76.98% at $n=0.134$, at redshift $z=0$ where the solutions differ the most - degree of accuracy the growth of structures in the context of $f(Q)$ power-law models. Consequently, this allowed us to rely on this approximation to analyze structure growth as presented in Sec. 4.
4 The structure growth in $f(Q)$-cosmology

### v4

@@ 3 Cosmological perturbations in $f(Q)$ gravity
where the first, second and last terms on the right hand side of the equation represent the frictional, restoring, and source forces, respectively, and the separation of variables takes the form $X=\sum_{k}X^{k}(t)Q^{k}(x),\hskip 2.84544pt{\rm and}\hskip 2.84544ptY=\sum_{k}Y^{k}(t)Q^{k}(x)\;,$ where $k$ is the wave number and $Q^{k}(x)$ is the eigenfunction of the covariantly defined Laplace-Beltrami operator in (almost) FLRW space-times, $\nabla^{2}Q^{k}(x)=-\frac{k^{2}}{a^{2}}Q^{k}(x).$ Then, the first-order evolution equations (28) - (31) become:
$\dot{\mathcal{D}}^{k}_{m}=-(1+w)Z^{k}+w\theta\mathcal{D}^{k}_{m}\;,$ (33)
$\dot{Z}^{k}=-\Big{[}\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}-\dot{Q}F^{\prime\prime}\Big{]}Z^{k}+\Bigg{[}\frac{w\theta^{2}}{3(1+w)}$
$-\frac{(1+3w)}{2(1+w)}\rho_{m}-\frac{wF}{2(1+w)}-\frac{QF^{\prime}w}{1+w}+\frac{\theta\dot{Q}F^{\prime\prime}w}{1+w}$
$+\frac{\theta\dot{Q}F^{\prime}w}{2Q(1+w)}+\frac{k^{2}}{a^{2}}\frac{w}{1+w}\Bigg{]}\mathcal{D}^{k}_{m}+\Bigg{[}\frac{1}{2}F^{\prime}-QF^{\prime\prime}$
$-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{]}\mathcal{W}^{k}$
$-\left[\frac{\theta F^{\prime}}{2Q}\mathcal{-}\theta F^{\prime\prime}\right]\mathcal{L}^{k}\;,$ (34)
$\dot{\mathcal{W}}^{k}=\mathcal{L}^{k}-\frac{w\dot{Q}}{1+w}\mathcal{D}^{k}_{m}\;,$ (35)
˙$\dot{\mathcal{L}}^{k}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}^{k}-\frac{w\ddot{Q}}{1+w}\mathcal{D}^{k}_{m}\;.$ (36)
@@
Consequently, the second-order evolution equations (33) - (36) with respect to time derivative yield as
$\ddot{\mathcal{D}}_{m}^{k}=-\left[\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}+\dot{Q}F^{\prime\prime}-w\theta\right]\dot{\mathcal{D}}_{m}^{k}-\left[-wF\right.$
$\left.-\theta\dot{Q}F^{\prime\prime}w+\frac{\theta\dot{Q}F^{\prime}w}{2Q}-\frac{(1+3w)\rho_{m}}{2}(1-w)\right.$
$\left.+\frac{k^{2}}{a^{2}}w\right]\mathcal{D}_{m}^{k}-\left[\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}\right.$
$\left.-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\right](1+w)\mathcal{W}^{k}$
$+\left[\frac{\theta F^{\prime}}{2Q}-\theta F^{\prime\prime}\right](1+w)\dot{\mathcal{W}}^{k}\;,$ (37)
$\ddot{\mathcal{W}}^{k}=\frac{\dddot{Q}}{\dot{Q}}\mathcal{W}^{k}-\frac{2w\ddot{Q}}{1+w}\mathcal{D}_{m}-\frac{w\dot{Q}}{1+w}\dot{\mathcal{D}}_{m}^{k}\,.$ (38)
3.1 Density contrast evolution
@@
$\dot{Y}=-(1+z)HY^{\prime}\;,$ (39)
$\ddot{Y}=(1+z)H^{2}Y^{\prime}+(1+z)^{2}H^{2}Y^{\prime\prime}+(1+z)^{2}H^{\prime}HY^{\prime}\;,$
@@
to present the growth density fluctuations through cosmological redshift. Thus, by resorting to expressions Eqs. (39) and assuming the power law model $F(Q)=\alpha Q^{n}$, the second-order evolution equations Eqs. (3) - (38) read as follows:
$(1+z)^{2}{\mathcal{D}}^{\prime\prime}_{m}=(1+z)\bigg{\{}1-(1+z)\frac{H^{\prime}}{H}+$
$\frac{H^{\prime}}{H}n\left({1-\Omega_{m}}\right)E^{2n-2}\bigg{\}}{\mathcal{D}}^{\prime}_{m}+\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\mathcal{D}_{m}$
$-\frac{1}{H^{2}}\Bigg{\{}\frac{(1-\Omega_{m})}{(1-2n)}E^{2n-2}\bigg{[}\frac{n^{2}}{2}-n-(1+z)\frac{{H^{\prime}}}{H}n\bigg{(}n^{2}$ (40)
$-4n+\frac{7}{2}\bigg{)}\bigg{]}\Bigg{\}}\mathcal{W}-(1+z)\frac{n}{2H^{2}}\left(1-\Omega_{m}\right)E^{2n-2}{\mathcal{W}}^{\prime}\;,$
$\left(1+z\right)^{2}\mathcal{W}^{\prime\prime}=\frac{40n(\Omega_{m}+3-\Omega_{m}-9)}{3(2n-1)(n^{2}\Omega_{m}-n^{2}+2n-1)}\mathcal{W}\;.~{}~{}~{}~{}~{}~{}~{}$ (41)
@@
As we can see, Eqs. (3) - (38) are nonlinear evolution equations in which no approximation has been made. To convert such a set of equations into a closed second-order equation, many works (Sahlu et al., 2020; Sami et al., 2021; Ntahompagaze et al., 2020; Abebe et al., 2013) have considered the quasi-static approximation. For simplicity, we applied the quasistatic approximation, where the first and second order of the time derivative of nonmetric density fluctuations is assumed to be approximately zero ($\dot{\mathcal{W}}=\ddot{\mathcal{W}}\approx 0$). Once the quasi-static approximation is made, Eqs. (3) - (38) reduced to the closed system equation and it becomes
$\ddot{\mathcal{D}}_{m}^{k}=-\Bigg{\{}\frac{2\theta}{3}+\frac{\dot{Q}F^{\prime}}{2Q}+\dot{Q}F^{\prime\prime}-w\theta+\Bigg{(}\frac{1}{2}F^{\prime}-QF^{\prime\prime}$
@@
$-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{)}\left(\frac{w\dot{Q}^{2}}{\dddot{Q}}\right)\Bigg{\}}\dot{\mathcal{D}}_{m}^{k}$
$-\Bigg{\{}\frac{\theta\dot{Q}F^{\prime}w}{2Q}-wF-\theta\dot{Q}F^{\prime\prime}w-\frac{(1+3w)\rho_{m}}{2}(1-w)$
$+\frac{k^{2}}{a^{2}}w+\Bigg{(}\frac{1}{2}F^{\prime}-QF^{\prime\prime}-F^{\prime}+\theta\dot{Q}F^{\prime\prime\prime}-\frac{\theta\dot{Q}F^{\prime\prime}}{2Q}$
@@
$+\frac{\theta\dot{Q}F^{\prime}}{2Q^{2}}\Bigg{)}\left(\frac{2w\dot{Q}\ddot{Q}}{\dddot{Q}}\right)\Bigg{\}}\mathcal{D}^{k}_{m}\,.$ (42)
For the case of $n=0$, Eqs. (42) reduce to the usual $\Lambda$CDM equation for the density contrast as
$\ddot{\mathcal{D}}_{m}^{k}=-\left(\frac{2}{3}-w\right)\theta\dot{\mathcal{D}}_{m}^{k}+\frac{1}{3}\theta^{2}\Big{[}2w(1-\Omega_{m})$
$+\frac{\Omega_{m}(1+3w)}{2}(1-w)-\frac{k^{2}}{3H^{2}a^{2}}w\Big{]}\mathcal{D}_{m}^{k}.$ (43)
Once redshift-space transformation is implemented, we find that quasi-static Eq. (42), for dust ($\omega=0$) and $f(Q)$ power-law models (12), yields
$(1+z)^{2}{\mathcal{D}}^{\prime\prime}_{m}=(1+z)\Bigg{[}1-(1+z)\frac{H^{\prime}}{H}+$ (44)
$\,\frac{H^{\prime}}{H}\left(1-\Omega_{m}\right)nE^{2n-2}\Bigg{]}{\mathcal{D}}^{\prime}_{m}+\,\frac{3\Omega_{m}}{2E^{2}}(1+z)^{3}\,\mathcal{D}_{m}\;.$
@@
To solve the density contrast numerically resorting the evolution equations (40)-(41), we use the initial values of the system. Due to gravitational instability, the variation in the temperature of the CMB is of the order $10^{-5}$ (Smoot et al., 1992) at $z_{in}\approx 1089$, and we consider it an initial condition. This small gravitational instability is the seed for the formation of large-scale structures. Hence, we define the normalized energy density contrast for the matter fluid as
$\delta(z)=\frac{\mathcal{D}_{m}(z)}{\mathcal{D}_{m}{(z_{in})}}\;,$ (45)
@@
where the subscript $in$ refers to the initial value of $\mathcal{D}_{m}(z)$ at the given initial redshift $z_{in}$ (Smoot et al., 1992). As stated previously, we consider two different situations in this section: the first illustration does not assume any (quasi-static) approximation (dubbed the full system in the following), whereas the second situation relies on the quasi-static approximation to determine the evolution of the density contrast $\delta_{m}(z)$ in the matter-dominated universe.
@@
Figure 2: The outer left panel displays growth density fluctuations, $\delta(z)$ when resolving the whole system (40)-(41), while the inner left panel illustrates the relative difference, $\zeta(z)$, between the $\Lambda$CDM and $f(Q)$-power-law models across the cosmological redshift. The outer right panel depicts growth density fluctuations $\delta(z)$ under the quasi-static approximation given by (3.1), and the inner right panel showcases the relative difference $\zeta(z)$, between the $\Lambda$CDM and $f(Q)$ power-law models. These plots correspond to values of $n=-0.085^{+0.300}_{-0.190}=\{0.215,-0.085,-0.275\}$ and $\Omega_{m}=0.304$, which are obtained from the best-fit parameters derived using OHD + SNIa data (see Table 1).
3.2 Numerical results of $\delta(z)$
We first consider the evolution equation without any approximation as presented in Eqs. (40) - (41), and show the results for the density contrast in the left outer panel of Fig. 2. using the best-fit values of $n$. Second, we have studied the growth of density fluctuations resorting to the quasi-static approximation as per Eq. (3.1) whose results are presented in the right outer panel of Fig. 2 for the matter-dominated universe. Meanwhile, in order to represent the density contrast $\delta(z)$ for the model $f(Q)$, for illustrative purposes, we have chosen the best-fit values of $\Omega_{m}=0.304$ and $n=-0.085^{+0.300}_{-0.190}=\{0.215,-0.085,-0.275\}$ as obtained from the analysis of OHD + SNIa presented in Table 1. For $\Lambda$CDM we have also chosen $\Omega_{m}=0.323$ so a comparison becomes possible.
@@ 3.2 Numerical results of $\delta(z)$
The numerical values of the fluctuation of matter density $\delta(z)$ turned out to be very sensitive to the constrained values of $\Omega_{m}$ and the exponent $n$ in both scenarios (full and quasi-static). We have also presented the relative difference of $\delta(z)$ between $\Lambda$CDM and $f(Q)$ model by defining the dimensionless parameter $\zeta(z)$ as
$\zeta(z)=\bigg{|}\frac{\delta^{\rm\Lambda CDM}(z)-\delta^{f(Q)}_{\rm{full,QSA}}(z)}{\delta^{\rm\Lambda CDM}(z)}\bigg{|}\;\;,$ (46)
@@
Figure 3: The relative difference between the full system and the quasi-static resolutions for $f(Q)$-gravity model for the values of $n=-0.085^{+0.300}_{-0.190}=\{0.215,-0.085,-0.275\}$and $\Omega_{m}=0.304$ which have been taken from the best-fit values obtained using OHD+SNIa data, see Table 1.
@@
$\eta(z)=\bigg{|}\frac{\delta^{f(Q)}_{\rm{QSA}}(z)-\delta^{f(Q)}_{\rm{full}}(z)}{\delta^{f(Q)}_{\rm{full}}(z)}\bigg{|}\;,$ (47)
and presented its numerical finding in Fig. 3. This definition serves as an indicator of the competitiveness of quasi-static with respect to the full system resolution. From Fig. 3, we conclude that the quasi-static approximation better represents the evolution of the density contrast for values of $n$ closer to zero ($n=0$ corresponding to $\Lambda$CDM). Based on this result, we highlight that the quasi-static approximation, which has been widely taken into consideration in other modified theories of gravity (Abebe et al., 2012; Sami et al., 2021; Abebe, 2015; Bose, 2018; Sawicki & Bellini, 2015) may be used to represent with a competitive 94.03% at $n=0.215$, 98.7% at $n=-0.085$, and 96.61% at $n=-0.275$, at redshift $z=0$ where the solutions differ the most, degree of accuracy, growth of structures in the context of $f(Q)$ power-law models. Consequently, this allowed us to rely on this approximation to analyse structure growth as presented in Sec. 4.
4 Structure growth in $f(Q)$ cosmology